a two digit number is 4 times to its sum of digits, when 9 is added to the number the digits will get reversed then what is that number

Grade:11

3 Answers

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

Lrt the 2 digit number be 10x+y

10x+y=4x+4y

6x=3y

2x=y

Now,

10x+y+9=10y+x

9x-9y=-9

x-y=-1

On solving we get

x=1

y=2

Number is : 10x+y=12

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

Sarita Sarkar
12 years ago

the number is 12 .

12 = 4* ( 1+2)

12 + 9 = 21, which is reverse the numberSmile

Sudheesh Singanamalla
12 years ago

2 digit number is 10x+y where x is tens place and y is units place

so

according to the question

10x + y = 4 (x+y)

solving it we get

6x-3y = 0 ------------------------(i)

the second part of the question says 9 is added to the number and then number obtained is the original numbers reverse

so

10x + y + 9 = 10y + x

solving we get

9x-9y=-9

==>  x - y = -1 --------------------------(ii)

 

multiplying (i) by 1 and  (ii) by 3 and subtracting (i) from (ii) we get

-3x = -3

==> x=1 ------------- (iii)

substituting this in any of the equations (i) or (ii) we can find y

y = 2

so the original number is 10x + y = 10*1 + 2 = 12

reversed number = 21 !

 

please approve if you like the answer